A sample of 0.255 mole of gas has a volume of 0.748L at 301K, Calculate the pressure of the gas.
(PV = nRT) (R=0.0821L x atm/ mol x K)
1
Expert's answer
2016-04-15T08:33:04-0400
PV = nRT, or
P = nRT/V

Therefore,
P = 0.255*0.0821*301/0.748 = 8.425 atm
Answer: 8.425 atm